OpenStudy (anonymous):

the length of a rectangle is four greater than the width,the area is sixty. find the length & with of the rectangle?

OpenStudy (anonymous):

let width =x , length =x+4 area = x(x+4) =60 =>x^2 + 4x -60 =0 => x^2 +10x -6x -60 = 0 => x(x +10) -6(x+10)=0 => (x-6)(x+10) = 0 => x=6, -10 as -10 is not possible... width = 6 unit, length = 10 unit